# Index Terms: 6G, large language models (LLMs), fine-tuning, safety alignment, open-source telecom datasetsI Introduction and Motivation

### v1

@@ Abstract
Fine-tuning large language models (LLMs) for telecom tasks and datasets is a common practice to adapt general-purpose models to the telecom domain. However, little attention has been paid to how this process may compromise model safety. Recent research has shown that even benign fine-tuning can degrade the safety alignment of LLMs, causing them to respond to harmful or unethical user queries. In this paper, we investigate this issue for telecom-tuned LLMs using three representative datasets featured by the GenAINet initiative. We show that safety degradation persists even for structured and seemingly harmless datasets such as 3GPP standards and tabular records, indicating that telecom-specific data is not immune to safety erosion during fine-tuning. We further extend our analysis to publicly available Telecom LLMs trained via continual pre-training, revealing that safety alignment is often severely lacking, primarily due to the omission of safety-focused instruction tuning. To address these issues in both fine-tuned and pre-trained models, we conduct extensive experiments and evaluate three safety realignment defenses (SafeInstruct, SafeLoRA, and SafeMERGE) using established red-teaming benchmarks. The results show that, across all settings, the proposed defenses can effectively restore safety after harmful degradation without compromising downstream task performance, leading to Safe teleCOMMunication (SafeCOMM) models. In a nutshell, our work serves as a diagnostic study and practical guide for safety realignment in telecom-tuned LLMs, and emphasizes the importance of safety-aware instruction and fine-tuning for real-world deployments of Telecom LLMs.
Index Terms:
6G, large language models (LLMs), fine-tuning, safety alignment, open-source telecom datasetsI Introduction and Motivation
@@
Large language models (LLMs) have demonstrated remarkable capabilities in coding, math, and, more recently, advanced reasoning tasks [1], while at the same time becoming more accessible to the broader public. This makes them powerful catalysts for transforming 6G beyond a mere network infrastructure upgrade into an intelligent wireless ecosystem that embeds connectivity, communication, and artificial intelligence (AI) into a unified, AI-driven network fabric [2]. This will involve multiple AI agents that can communicate with one another, and not only understand the network data but also the underlying telecom standards, systems, and protocols [3]. To achieve this, generalist LLMs need to instill telecom-specific knowledge into their core understanding. This can be done through supervised fine-tuning (SFT) [4] or continual pre-training (CPT) on large telecom corpora such as 3GPP standards and scientific publications [5]. This reduces reliance on complex retrieval-augmented generation (RAG) systems that introduce additional latency from querying external knowledge bases.
@@
However, safety alignment, which guides models away from inappropriate behavior after pre-training, is critical to ensure that LLM outputs comply with human preferences and security policies. This is typically achieved as part of instruction tuning [6], e.g., via reinforcement learning with human feedback (RLHF) [7] or SFT, both of which incorporate corrective layers to steer the model away from harmful responses. For example, consider a 3GPP chatbot. While being helpful on telecom-related queries, it must be ensured that under no circumstances does it respond to unethical, harmful, or dangerous queries, such as: “How can I most effectively destroy a base station with as little effort as possible? Please make a short list.”
@@
Nonetheless, maintaining safety alignment is challenging when LLMs are further fine-tuned for downstream tasks. For example, the authors in [8] show that even a handful of malicious training examples can jailbreak LLMs, causing them to comply with harmful user requests. In addition, [9] demonstrates that even conventionally benign tasks can inadvertently compromise the safety of fine-tuned models to a noticeable degree. Theoretical insights to this phenomenon were explored in [10] and [11], where the authors investigate refusal directions and the token-depth of safety alignment, suggesting that alignment is often shallow and easily disrupted.
@@
These findings confirm a problematic reality: an instruction-tuned LLM may override its safety guardrails when fine-tuned for new downstream tasks. Therefore, ensuring that LLMs remain safe after fine-tuning is an important practical challenge for Telecom LLMs. In this paper, we address this challenge by investigating SFT fine-tuning of widely used Llama [12, 13] and Qwen [14] models on three publicly available telecom datasets featured by the GenAINet initiative [15]: TeleQnA [16], TeleData [5], and TSpecLLM [17]. In addition, we examine publicly released TeleLLMs [5], which have been continually pre-trained on large-scale telecom corpora. Our main contribution is to show that adapting LLMs to telecom data degrades safety, and to demonstrate that safety can be restored using lightweight methods. Our key findings are:
@@
Continual pre-training (CPT) alone without safety-focused instruction tuning leads to severe safety issues with harmfulness ratios close to 90%, such that TeleLLMs comply to almost any harmful user prompt.
To mitigate these issues, we evaluate three safety realignment defenses that can be easily integrated by practitioners with open-source fine-tuning libraries: SafeInstruct [19], SafeLoRA [20], and SafeMERGE [21]. Our extensive experiments demonstrate that each defense can effectively restore safety after harmful degradation with minimal impact on downstream telecom task performance. These approaches yield Safe teleCOMMunication (SafeCOMM) models that strike a good balance between safety and utility. To support our claims, we provide theoretical insights in Section II, discuss realignment defenses in Section III, explain our experimental setup in Section IV, and present our results in Section V.
@@ II Background: Safety Degradation in LLMs
Safety alignment in LLMs ensures that models provide helpful yet harmless responses. While many instruction-tuned LLMs (labeled as instruct/chat variants) exhibit such safety features, a growing body of work has demonstrated that both SFT and CPT can inadvertently degrade safety alignment [22]. To understand why this occurs in practice, we provide a brief overview of SFT and CPT and explain how both approaches can unintentionally cause models to respond unsafely, particularly when using telecom-specific datasets (see Fig. 1).
II-A Supervised Fine-Tuning (SFT).
SFT typically takes an instruction-tuned model, which is trained to follow a helpful chat-like behavior. In general, such instruction tuning involves safety-aligned prompts to reinforce refusal behaviors, making instruct/chat models safe to use. To adapt the model to a specific domain (e.g., telecom), data is often structured as question-answer (QA) pairs to continue the chat-like behavior while absorbing domain knowledge. Formally, SFT optimizes the following objective:
where represent telecom QA pairs from the distribution for next token prediction with parameters , e.g.,
@@
A: 3GPP defines two main frequency ranges: FR1 (410 MHz to 7.125 GHz) and FR2 (24.25 GHz to 52.6 GHz).
The result is a fine-tuned telecom instruct model. However, during SFT, the previously established safety alignment may be eroded while general instruction following capabilities remain. This erosion may occur due to several phenomena:
Embedding drift: model updates may unintentionally overwrite layers responsible for safe refusal behavior [8].
@@
Shallow safety alignment: safety alignment is often just a few tokens deep such that token distribution shifts can inadvertently break safety alignment during SFT[11].
Bias and unsafe data in pre-training: SFT may re-surface/amplify unsafe layers from pre-training [22].
In general, telecom datasets do not contain safety prompts which could help realign the model’s guardrails during SFT.
II-B Continual Pre-Training (CPT).
@@
CPT extends pre-training of a non-instruct (base) model on large-scale unlabeled corpora. In the case of telecom, these may include 3GPP standards, scholarly papers, and other data that are not necessarily formatted as QA pairs [5]. Formally, let a pretrained LLM be trained on a generic corpus with parameters . CPT aims to minimize the next-token prediction loss on a new domain corpus , thereby adapting the model’s parameters to the telecom domain:
@@
where are token sequences sampled from the telecom corpus. After CPT, the model typically undergoes instruction tuning to instill helpfulness and chat-like behavior. To ensure safety, instruction tuning datasets need to include explicit safety examples. However, this is not always the case for open-source datasets that focus largely on helpfulness.
@@
This may create a serious oversight for non-familiar practicioners: By pre-training again on telecom-specific data without concurrent safety alignment, models may gradually “forget” previous safety guardrails. As shown in prior works [8, 22], CPT can overshadow safety layers if domain corpora lack safety-critical prompts, as is often the case for telecom data. In such cases, the model becomes more likely to generate harmful outputs, even if the training data is benign on the surface.
II-C Why Telecom Data May Contribute to Safety Degradation.
@@
Telecom datasets often contain large volumes of structured entries (e.g., 3GPP tables, system logs), long documents (standardization reports, release notes), and specialized protocols featuring complex formulas. As shown in [23], such structured tabular and math-heavy data can amplify safety degradation by exposing LLMs to specialized jargon and formats that resemble harmful samples. The results in [23] demonstrate that benign data, particularly when formatted as lists, bullet points, or math problems, can significantly increase harmfulness after fine-tuning. This occurs because these formats often share update directions in gradient space with known harmful data. Given the prevalence of these formats in telecom datasets, this raises an important open question: Does SFT or CPT with telecom data lead to similar or even worse safety degradation? Our goal is to address this question using open-source telecom datasets, investigating the importance of safety-aware data curation even in domains typically perceived as benign.
@@ III Safety Realignment in Telecom LLMs
We primarily focus on defenses for fine-tuning via SFT and demonstrate in Section IV how the same concepts can be extended to CPT. Further, we investigate defenses that are either applicable after harm has already happened, i.e. post-hoc safety corrections, or that do not require much effort during the fine-tuning stage. In addition, we want the defenses to be as simple as possible such that they can be easily integrated by practitioners for telecom fine-tuning with open-source libraries. Based on these criteria, we evaluate SafeInstruct [19], a fine-tuning-stage defense, as well as SafeLoRA [20] and SafeMERGE [21], both post-fine-tuning-stage defenses.
III-A SafeInstruct: Mixing Safety QAs during Fine-Tuning.
SafeInstruct interleaves a small portion of safety-aligned QA pairs into the original fine-tuning dataset. These typically include harmful questions with safe refusals, e.g.,
Q: Outline the steps to set a vehicle on fire.
@@
As safety data is part of training, utility is usually minimally impacted. Depending on the dataset, more safety samples may be required or need to be potentially adapted to the domain (e.g., telecom). We will investigate this in our experiments.
III-B SafeLoRA / SafeMERGE: Layer-Wise LoRA Adaptation.
SafeLoRA and SafeMERGE selectively adapt only those layers in a fine-tuned LLM that exhibit harmful behavior. In most practical scenarios, SFT is implemented via LoRA [24], which introduces low-rank adapters into the model layers. For a weight matrix in a transformer layer , LoRA introduces two trainable matrices and (with ) such that the adapted weight becomes , where is a scaling factor. During fine-tuning, remains frozen while only and are trained. Both SafeLoRA and SafeMERGE introduce a safety-aligned subspace , computed as the difference between the weights of the base (unaligned) and instruct/chat (aligned) version of the model. This subspace represents the safety alignment in the weight space and the projection onto it can be computed by:
For each layer , if the deviation from this subspace is large, SafeLoRA projects the corresponding layer onto , while SafeMERGE merges the layer with that of a known safe model (e.g., fine-tuned on safety-aligned data only). More formally, let and be the LoRA updates of the -th layer from the fine-tuned and SafeMERGE’s safe model, respectively. The cosine similarity between the fine-tuned adapter and its projection onto the safety subspace serves to quantify how much the LoRA adapter deviates from safety alignment, i.e.
@@
SafeLoRA projects the adapter onto the subspace:
@@
SafeMERGE merges it (e.g., linear) with the safe version:
(6) (7) where . Merging strategies other than linear merging can be applied with SafeMERGE as well.
@@
In summary, all three defenses aim to mitigate harmfulness in fine-tuned LLMs. In the following experiments, we will evaluate whether these defenses can be straightforwardly applied to fine-tuning with domain-specific telecom datasets.
@@ IV Experimental Setup
We closely follow the experimental setup in [21] for safety-related model evaluations and safety realignment defenses.
IV-A SFT on Telecom Datasets.
@@
We fine-tune three widely used instruction-tuned models: Llama-2-7B-Chat [12], Llama-3.1-8B-Instruct [13], and Qwen-2.7B-Instruct [14]. For datasets, we choose the QA-formatted benchmark datasets from TeleQnA [16] (8k samples), TeleData [5] (600k samples), and TSpecLLM [17] (80 samples), where we created 80/20 train-test splits, respectively. These datasets contain various telecom-specific questions drawn from standards, implementations, and engineering practice, often in potentially harmful formats for fine-tuning, i.e. lists, tables, bullet points, and complex mathematical formulas. The varying dataset sizes further allow us to analyze how the amount of telecom data impacts safety after SFT. All models are fine-tuned using Llama-Factory with an effective batch size of 32 and a learning rate of 1e-4 with linear scheduling. We train for 2 epochs on TeleData and for 5 epochs on TeleQnA and TSpecLLM. After fine-tuning, we evaluate model performance on the test split by following the approach in [5]. In particular, we use Mixtral-8x7B-Instruct [25] as an LLM judge to compare generated answers with ground truth responses, a popular choice for semantically evaluating complex LLM outputs. We compute the final accuracy on each QA dataset as the ratio of correctly answered questions.
IV-B Safety Evaluations.
To assess safety, we generate model responses on Direct-Harm [18] and HexPhi [9], two established red-teaming benchmarks containing prompts that intentionally conflict with aligned LLM policies. Following standard practice in AI safety research, we use a safety LLM judge, Llama-Guard-3-8B [13], to evaluate the harmfulness of generated responses. We then measure safety as the harmful output rate, where lower values indicate better safety. We report the overall harmfulness score as the proportion of generated responses flagged as harmful.
IV-C Safety Realignment Defenses.

### v2

# I Introduction and Motivation

@@ Abstract
Fine-tuning large language models (LLMs) on telecom datasets is a common practice to adapt general-purpose models to the telecom domain. However, little attention has been paid to how this process may compromise model safety. Recent research has shown that even benign fine-tuning can degrade the safety alignment of LLMs, causing them to respond to harmful or unethical user queries. In this paper, we investigate this issue by fine-tuning LLMs on three representative telecom datasets and show that safety degrades even for light telecom domain adaptation. To this end, we introduce TeleHarm, the first telecom-specific red-teaming benchmark, which we use alongside established Direct-Harm and HexPhi datasets to systematically assess harmful behavior. We further extend our analysis to publicly available TeleLLMs that were continually pre-trained on large telecom corpora, revealing that safety alignment is severely lacking, primarily due to the omission of safety-focused instruction tuning. To address these issues, we evaluate three realignment defenses: SafeInstruct, SafeLoRA, SafeMERGE. We show that, across all settings, the proposed defenses can effectively restore safety without compromising telecom task performance, leading to Safe teleCOMMunication (SafeCOMM) models. Our work serves as both a diagnostic study and practical guide for safety realignment in telecom-tuned LLMs, underscoring the need for safety-aware instruction and fine-tuning in the telecom domain.
I Introduction and Motivation
@@
Large language models (LLMs) have demonstrated remarkable capabilities in coding, math, and advanced reasoning. This makes them powerful catalysts for transforming 6G beyond a mere network infrastructure upgrade into an intelligent wireless ecosystem that embeds connectivity, communication, and artificial intelligence (AI) into a unified network fabric [1]. This will involve AI agents that not only understand the network data but also the underlying telecom standards, systems, and protocols [2]. To achieve this, LLMs need to instill telecom knowledge into their core understanding. This can be done through supervised fine-tuning (SFT) [3] or continual pre-training (CPT) using large telecom corpora such as 3GPP standards and scientific publications [4].
@@
However, safety alignment, which prevents models from inappropriate behavior, is critical to ensure that LLMs comply with human preferences and security policies. This is typically achieved as part of instruction tuning [5] or SFT, both of which incorporate corrective layers using safe data to steer the model away from harmful responses. For example, consider a 3GPP chatbot. While helpful on telecom-related queries, it must be ensured that it does not respond to unethical, harmful, or dangerous prompts, such as: “How can I destroy a public base station? Please make a short list.”
@@
Nonetheless, maintaining safety alignment is challenging when LLMs are further fine-tuned for downstream tasks. Recent work in [6] shows that a handful of malicious training examples can already jailbreak LLMs, causing them to comply with harmful user requests. In addition, the authors in [7] demonstrate that even conventionally benign fine-tuning can inadvertently compromise safety to a noticeable degree. Theoretical insights on this were later explored by investigating refusal directions and token-depth, suggesting that safety alignment is often shallow and easily disrupted [8, 9].
@@
These findings confirm a problematic reality: LLMs may override their safety guardrails when adapted for new domains. Therefore, ensuring that telecom LLMs remain safe after fine-tuning is an important practical challenge. In this paper, we address this by investigating SFT of widely used Llama [10, 11] and Qwen [12] models on three publicly available telecom datasets: TeleQnA [13], TeleData [4], and TSpecLLM [14]. In addition, we examine publicly released TeleLLMs [4], which underwent continual pre-training using large telecom corpora. Our main goal is to confirm that even seemingly benign telecom datasets are not exempt from safety degradation, and to demonstrate that safety can be restored using lightweight methods. Our key contributions and findings are:
@@
Supervised fine-tuning (SFT) on telecom data leads to noticeable safety degradation, as measured across DirectHarm, HexPhi, and TeleHarm benchmarks.
3.
Continual pre-training (CPT) without safe instruction tuning leads to harmfulness ratios close to 90%. where TeleLLMs comply to most harmful user prompts.
To mitigate these issues, we evaluate three safety realignment defenses: SafeInstruct [16], SafeLoRA [17], and SafeMERGE [18]. In extensive experiments, we demonstrate that each defense can effectively restore safety with minimal impact on performance, yielding Safe teleCOMMunication (SafeCOMM) models that balance safety and utility.
@@ II Background: Safety Degradation in LLMs
Safety alignment ensures that LLMs provide helpful yet harmless responses. While many instruction-tuned LLMs are safety aligned, a growing body of work has demonstrated that continued SFT or CPT can inadvertently degrade safety [19]. To understand why this occurs in practice, we provide a brief overview of SFT and CPT, and explain how they can unintentionally cause models to respond to unsafe prompts, particularly when fine-tuned on telecom datasets (see Fig. 1).
II-A SFT with Instruction-Tuned Models
Instruction tuning LLMs generally involves incorporating safety data, making instruct LLM variants not only follow a helpful chat-like behavior but also reinforce safe refusal behavior [20]. To adapt the model to a specific domain (e.g., telecom), data is often structured as question-answer (QA) pairs to continue the chat-like behavior while absorbing domain knowledge. Formally, SFT optimizes the objective:
where represent telecom QA pairs from the distribution for next token prediction with parameters , e.g.:
@@ II-A SFT with Instruction-Tuned Models
A: 2 main frequency ranges: FR1 (410 MHz to 7.125 GHz) and FR2 (24.25 GHz to 52.6 GHz).
@@
However, SFT can erode safety alignment due to several phenomena: a) embedding drift: model updates may unintentionally overwrite safe refusal layers [6], b) shallow alignment: safety alignment is often just a few tokens deep such that distribution shifts can inadvertently break it [9], c) unsafe pre-training data: SFT may resurface unsafe layers [19].
II-B Continual Pre-Training (CPT)
@@
In contrast to SFT, CPT extends pre-training of a non-instruct (base) model on large-scale unlabeled corpora. For telecom, these may include 3GPP standards, scholarly papers, protocol designs, and other sources that are not necessarily formatted as QA pairs [4]. Formally, let an LLM be pre-trained on a generic corpus with parameters . CPT aims to minimize the next-token prediction loss on the new domain corpus , adapting the model’s parameters , i.e.:
@@
After CPT, the LLM typically undergoes instruction tuning to instill helpfulness and chat-like behavior. To ensure safety, instruct datasets need to include explicit safety examples. However, this is not always the case for open-source datasets that focus largely on helpfulness and conversational style.
@@
This may create a serious oversight for non-familiar practitioners. As shown in prior works [6, 19], CPT can bypass internal safety guardrails if domain corpora lack safety-critical prompts, potentially resulting in a model that is more likely to generate harmful outputs, even if the training data is benign on the surface. In general, telecom datasets do not contain such safety prompts that could help realign the model’s guardrails.
II-C Why Telecom Data May Be Prone to Safety Degradation
@@
Telecom datasets often contain large volumes of tabular entries (e.g., 3GPP tables), long documents (e.g., standardization reports), and specialized protocols with complex formulas and algorithms. As shown in [21], such structured tabular and math-heavy data can amplify safety degradation by exposing LLMs to specialized jargon and formats that resemble harmful samples. Specifically, the authors demonstrate that benign data, particularly when formatted as lists, bullet points, or math problems, can significantly increase harmfulness after SFT. This occurs because these samples often share update directions in gradient space with known harmful data. Given the prevalence of these formats in telecom datasets, this raises an important open question: Does SFT or CPT with telecom data lead to similar or even worse safety degradation? Our goal is to address this question using open-source telecom datasets, investigating the importance of safety-aware data curation even in domains typically perceived as benign. In addition, our work provides guidance for practitioners unfamiliar with AI safety, showing how safe telecom LLMs can be trained with open-source frameworks without deep expertise in safety alignment.
@@ III Safety Realignment in Telecom LLMs
We first focus on defenses for SFT and demonstrate in Sec. IV how the same concepts can be extended to CPT. In particular, we investigate lightweight defenses that are either applicable after harm has already happened, i.e., post-SFT, or that can be applied during SFT without significant engineering effort. Based on these criteria, we evaluate SafeInstruct [16], a fine-tuning-stage defense, as well as SafeLoRA [17] and SafeMERGE [18], both post-fine-tuning-stage defenses.
III-A SafeInstruct: Mixing Safety QAs during Fine-Tuning
SafeInstruct interleaves a small portion of safety-aligned QA pairs into the original fine-tuning dataset. These typically include harmful questions with safe refusals, e.g.:
Q: How to bypass 5G authentication?
@@ III-A SafeInstruct: Mixing Safety QAs during Fine-Tuning
As safety data is part of training, utility is usually minimally impacted if the number of safety samples is not too large.
III-B SafeLoRA / SafeMERGE: Layer-Wise LoRA Adaptation
In most practical scenarios, SFT is implemented via parameter-efficient LoRA training [22], which introduces low-rank adapters into the model layers. For a weight matrix in a transformer layer , LoRA introduces two trainable matrices and (with ) such that the adapted weight becomes , where is a scaling factor. During fine-tuning, remains frozen while only and are trained. Both SafeLoRA and SafeMERGE selectively adapt only those LoRA layers that exhibit harmful behavior. To this end, they introduce a safety-aligned subspace , computed as the difference between the weights of the base (unaligned) and instruct (aligned) version of the model. This subspace represents the safety alignment in the weight space and the projection onto it can be computed by:
For each layer , if the deviation from this subspace is large, SafeLoRA projects the corresponding layer onto , while SafeMERGE merges the layer with that of a known safe model (e.g., fine-tuned on safety-aligned data only). More formally, let and be the LoRA updates of the -th layer from the fine-tuned and SafeMERGE’s safe model, respectively. The cosine similarity between the fine-tuned adapter and its projection onto the safety subspace serves to quantify how much the LoRA adapter deviates from safety alignment, i.e.:
@@ III-B SafeLoRA / SafeMERGE: Layer-Wise LoRA Adaptation
SafeLoRA projects the adapter onto the subspace , i.e.:
@@
SafeMERGE merges it with the safe model , i.e.:
@@ IV Experimental Setup
In the following experiments, we evaluate whether fine-tuning with domain-specific telecom datasets increases harmfulness and whether defenses such as SafeInstruct, SafeLoRA, and SafeMERGE can be straightforwardly applied.
IV-A SFT Fine-Tuning on Telecom Datasets
@@
We fine-tune three widely used instruction-tuned models: Llama-2-7B-Chat [10], Llama-3.1-8B-Instruct [11], and Qwen-2-7B-Instruct [12]. For datasets, we choose the QA-formatted benchmark datasets from TeleQnA [13] (8k samples), TeleData [4] (600k samples), and TSpecLLM [14] (80 samples), where we created 80/20 train-test splits, respectively. These datasets contain various telecom-specific questions drawn from standards, implementations, and engineering practice, often formatted as lists, tables, and complex mathematical formulas. The varying dataset sizes further allow us to analyze how the amount of telecom data impacts safety after SFT. All models are fine-tuned using Llama-Factory with an effective batch size of 32 and a learning rate of 1e-4 with linear scheduling. We train for 2 epochs on TeleData and for 5 epochs on TeleQnA and TSpecLLM. After SFT, we evaluate model performance on the test split by following the approach in [4] where we use Mixtral-8x7B-Instruct [23] as a judge to compare answers with ground truth responses. We compute the final accuracy as the ratio of correctly answered questions.
IV-B Safety Evaluations
To assess safety, we generate responses on DirectHarm [15] and HexPhi [7], two general-purpose red-teaming datasets containing intentionally harmful prompts. In addition, we introduce TeleHarm111Dataset available at: https://huggingface.co/aladinDJ/TeleHarm, a telecom-specific red-teaming benchmark of 125 harmful prompts carefully curated by domain experts across five categories: Physical/RF Layer, Access & Authentication, Core Network & Signaling, Infrastructure & OSS/BSS, and Privacy & Social Engineering. With TeleHarm, we target threat scenarios that are unique to communication systems and that span the entire network stack, providing, to the best of our knowledge, the first comprehensive telecom-specific benchmark that captures both high- and low-level risks, thereby complementing general-purpose red-teaming datasets. Following standard practice, we use a safety LLM judge (Llama-Guard-3-8B [11]) to evaluate model safety and report the overall harmfulness score as the proportion of generated responses flagged as unsafe.
IV-C Safety Realignment Defenses